OpenStudy (anonymous):

(3x)(x^3)^8 simplify please

OpenStudy (campbell_st):

1st step use the power of a power rule \[(x^a)^b = x^{a \times b}\] 2nd step when multiplying the same base add the powers \[x^a \times x^b = x^{a +b}\]

OpenStudy (anonymous):

I'm assuming it is \[3x(x^3)^8\] You can start with the left term, and remember that with an exponent on an exponent, you multiply them. So you end up with 3x(x^24). Then using the same principal as \[x*x^2=x^3\] it is just 3x^25.

OpenStudy (anonymous):

so is 3x^25 is that the answer

OpenStudy (campbell_st):

yes

OpenStudy (anonymous):

thank you